OpenStudy (anonymous):

factor the trinomial completely 12x^3-27x^2-27x

OpenStudy (tkhunny):

Hunt for common factors. 12x^3 = 2*2*3*x*x*x 27x^2 = 3*3*3*x*x 27x = 3*3*3*x What's common?

OpenStudy (anonymous):

2,3,x

OpenStudy (tkhunny):

2? No. That appears in only one term. There is no common 2. There is a common 3. There is a common x. 3x is the common factor that will help us. The Distributive Property suggests: 12x^3-27x^2-27x = 3x(4x^2 - 9x - 9)

OpenStudy (anonymous):